OpenStudy (anonymous):

Solve the system y = x + 2 and y = -x - 3

OpenStudy (anonymous):

If you know what y equals (from the first equation), you can substitute it into the second equation and solve for x. x+2=-x-3 (after substituting x+2 for y) 2x=-5 x=-2.5 Then substitute in the value for x that you found into the first equation to find y. y=-2.5+2 y=-0.5 You can then check your answer by substituting the values for x and y into both equations to make sure they work out properly.

OpenStudy (anonymous):

Can you help with this? 2x - 3y = 7 and 8x - 12y = 5

OpenStudy (anonymous):

My guess is that the second system is unsolvable.

OpenStudy (anonymous):

If you multiply both sides of the first equation by 4, you get 8x-12y=28. 8x-12y can't be both 28 and 5.